OpenStudy (anonymous):

solve the solution by the elimination method 5x+3y=-13 7x-2y=17

OpenStudy (anonymous):

sorry,solve the system

OpenStudy (anonymous):

wanna cheat?

OpenStudy (anonymous):

there are no cheats in math, only more clever methods

OpenStudy (anonymous):

Solve for y for both equations. Set y = y. Then solve for x.

OpenStudy (anonymous):

answer is \[x=\frac{25}{31},y=-\frac{176}{31}\]

OpenStudy (anonymous):

cheating is learning

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

No it isn't, lol

OpenStudy (anonymous):

it is how i learned for sure. we just called it "research"

OpenStudy (anonymous):

Hmmm

OpenStudy (anonymous):

http://www.youtube.com/watch?v=IL4vWJbwmqM

OpenStudy (anonymous):

so im guessing the answer isnt coming from satellite

OpenStudy (anonymous):

i sent it already

OpenStudy (anonymous):

5x + 3y = - 13 5x = -13-3y x= (-13-3y)/5 7(-13-4y)/5 - 2y = 7 (-91-28y)/5 - 2y = 7 (-91-28y-10y)/5 = 7 (-91-38y) = 35 -38y = 35 + 91 -38y = 126 y = 126/-38 y = -126/38 y = -63/19 5x + 3(-63/19) = -13 5x -189/19 =- 13 (95x - 189)/19 = -13 95x - 189 = -247 95x = -247 + 189 95x = -58 x = -58/95

OpenStudy (anonymous):

here it is again. \[x=\frac{25}{31},y=-\frac{176}{31}\]

OpenStudy (anonymous):

ordered pair

OpenStudy (anonymous):

c'mon. \[(\frac{25}{31},-\frac{176}{31})\]

OpenStudy (anonymous):

I'm getting x = 25/31

OpenStudy (anonymous):

got it or not cali

OpenStudy (anonymous):

I don't know how scientist got his answer

OpenStudy (anonymous):

ok one more time \[(\frac{25}{31},-\frac{176}{31})\]

OpenStudy (anonymous):

there may be some calculations mistake

OpenStudy (anonymous):

I think Cali may have been waiting for more than one person to post the same answer

OpenStudy (anonymous):

5x+3y=-137_x-2y=17 Since -2y does not contain the variable to solve for, move it to the right-hand side of the equation by adding 2y to both sides. 5x+3y=-137_x=2y+17 Replace all occurrences of x with the solution found by solving the last equation for x. In this case, the value substituted is 2y+17. 5(2y+17)+3y=-137_x=2y+17 Multiply 5 by each term inside the parentheses. (10y+85)+3y=-137_x=2y+17 Since 10y and 3y are like terms, add 3y to 10y to get 13y. 13y+85=-137_x=2y+17 Since 85 does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 85 from both sides. 13y=-85-137_x=2y+17 Subtract 137 from -85 to get -222. 13y=-222_x=2y+17 Divide each term in the equation by 13. (13y)/(13)=-(222)/(13)_x=2y+17 Simplify the left-hand side of the equation by canceling the common terms. y=-(222)/(13)_x=2y+17 Replace all occurrences of y with the solution found by solving the last equation for y. In this case, the value substituted is -(222)/(13). y=-(222)/(13)_x=2(-(222)/(13))+17 Multiply 2 by each term inside the parentheses. y=-(222)/(13)_x=-(444)/(13)+17 To add fractions, the denominators must be equal. The denominators can be made equal by finding the least common denominator (LCD). In this case, the LCD is 13. Next, multiply each fraction by a factor of 1 that will create the LCD in each of the fractions. y=-(222)/(13)_x=17*(13)/(13)-(444)/(13) Complete the multiplication to produce a denominator of 13 in each expression. y=-(222)/(13)_x=(221)/(13)-(444)/(13) Combine the numerators of all fractions that have common denominators. y=-(222)/(13)_x=(221-444)/(13) Subtract 444 from 221 to get -223. y=-(222)/(13)_x=(-223)/(13) Move the minus sign from the numerator to the front of the expression. y=-(222)/(13)_x=-(223)/(13) This is the solution to the system of equations. y=-(222)/(13)_x=-(223)/(13)

OpenStudy (anonymous):

dubwise http://www.youtube.com/watch?v=1S9078O6Il8

OpenStudy (anonymous):

Scientist!!!

OpenStudy (anonymous):

I think your input was wrong

OpenStudy (anonymous):

What is 7_x ?

OpenStudy (anonymous):

heavyweight dub champion. dub barrington levy robin hood all time greatest

OpenStudy (anonymous):

Scientist...Bagratrix is a fraud

OpenStudy (anonymous):

Always give wrong answers

OpenStudy (anonymous):

bliss it is not no bliss

OpenStudy (anonymous):

it never gives wrong answers

OpenStudy (anonymous):

Hold on, let me check...

OpenStudy (anonymous):

Yep, it's wrong...x is supposed to be positive, not negative

OpenStudy (anonymous):

Look at the second equation...If x is negative, no way you get positive 17 for the answer....sorry

OpenStudy (anonymous):

it solved with substitution method

OpenStudy (anonymous):

Besides, x is supposed to be a value less than 1. You have something much greater than 1

OpenStudy (anonymous):

5x+3y=-13_7x-2y=17 Multiply each equation by the value that makes the coefficients of y equal. This value is found by dividing the least common multiple of the coefficients of y by the current coefficient. In this case, the least common multiple is 6. 2*(5x+3y=-13)_3*(7x-2y=17) Multiply each equation by the value that makes the coefficients of y equal. This value is found by dividing the least common multiple of the coefficients of y by the current coefficient. In this case, the least common multiple is 6. 2*(5x+3y)=2(-13)_3*(7x-2y)=3(17) Multiply 2 by each term inside the parentheses. 2*(5x+3y)=-26_3*(7x-2y)=3(17) Multiply 2 by each term inside the parentheses. (10x+6y)=-26_3*(7x-2y)=3(17) Remove the parentheses around the expression 10x+6y. 10x+6y=-26_3*(7x-2y)=3(17) Multiply 3 by each term inside the parentheses. 10x+6y=-26_3*(7x-2y)=51 Multiply 3 by each term inside the parentheses. 10x+6y=-26_(21x-6y)=51 Remove the parentheses around the expression 21x-6y. 10x+6y=-26_21x-6y=51 Add the two equations together to eliminate y from the system. 21x-6y=51_<U>10x+6y=-26<u>_31x =25 Divide each term in the equation by 31. x=(25)/(31) Substitute the value found for x into the original equation to solve for y. 10((25)/(31))+6y=-26 Multiply 10 by each term inside the parentheses. (250)/(31)+6y=-26 Move all terms not containing y to the right-hand side of the equation. 6y=-(1056)/(31) Divide each term in the equation by 6. y=-(176)/(31) This is the final solution to the independent system of equations. x=(25)/(31)_y=-(176)/(31)

OpenStudy (anonymous):

Oh wow, now your answers are magically correct...

OpenStudy (anonymous):

got it now bliss it is never wrong before i did it with substitution method but now i did it with addition/elimination method

OpenStudy (anonymous):

There's only one way to get it right everytime

OpenStudy (anonymous):

good lord what a lot of work.

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

yes very much

OpenStudy (anonymous):

Sometimes a lot of hard work is wasted if the answer is wrong

OpenStudy (anonymous):

yes but no problem